# Fast Ewald Summation using Prolate Spheroidal Wave Functions

### v1

@@ 6.2.1 Choosing $c_{s}$
A simple choice for $c_{s}$ is to use the asymptotic estimate $c_{s}\sim\log(1/\varepsilon)$ [17, 19]. However, this choice tends to overestimate $c_{s}$ for the small to moderate values of $c_{s}$ (see Figure 4). Using the fitted model (91), we instead recommend choosing $c_{s}$ as
$c_{s}\approx\frac{1}{2}W\Big(2\,\big(B_{s}/\varepsilon\big)^{2}\Big),\qquad B_{s}:=\frac{1}{\sqrt{V}}A_{s}\|\bm{\rho}\|_{2}\sqrt{r_{c}}.$ (120)
where $W$ is the Lambert $W$ function.
@@ 6.3.1 Choosing $c_{w}$
To choose $c_{w}$, we assume that the split and window errors are equal. Balancing (115) and (91) we get
@@
We then solve (125) for $c_{w}$ using the Lambert $W$ function and obtain
$c_{w}\approx-\frac{1}{2}W\bigg(\!-2B_{w}^{2}\frac{e^{-2c_{s}}}{c_{s}}\bigg),\qquad B_{w}:=\frac{A_{s}}{A_{w}}\sqrt{\frac{r_{c}}{L}}.$ (126)
@@ 6.4 Parameter selection procedure
1: Solve $c_{s}=\frac{1}{2}W\big(2\,(B_{s}/\varepsilon)^{2}\big)$ for $c_{s}$, where $B_{s}:=\frac{1}{\sqrt{V}}A_{s}\|\bm{\rho}\|_{2}\sqrt{r_{c}}$.
@@ 7.1 Experimental setup
Unless otherwise stated, we consider systems of $n=100$ particles uniformly distributed in $0,1)^{3}$, generated using MATLAB’s rand function. The particle charges $\rho_{j}$ are sampled independently from a standard normal distribution (mean 0, variance $1$) and subsequently shifted to enforce charge neutrality, $\sum_{j}\rho_{j}=0$. The choice $n=100$ is sufficient to capture the relevant convergence trends; larger systems exhibit the same slopes and offsets; see e.g. Figure 7. Unless otherwise stated, we set the cutoff radius to $r_{c}=0.1$ and the domain size to $L=1$, which provides reasonable Fourier-space resolution for testing purposes.
We compute the approximate far-field potential values $\bm{\phi}^{\mathrm{far}}=(\phi^{\mathrm{far}}_{j})_{j=1}^{n}$ using Algorithm 2 and compare them with a reference solution $\bm{\phi}^{\mathrm{far}}_{\mathrm{ref}}=(\phi^{\mathrm{far}}_{\mathrm{ref},j})_{j=1}^{n}$, obtained by direct Ewald summation (Algorithm 1). In the reference calculation, we employ a sufficiently large number of Fourier modes so that the residual error is at machine precision. The Fourier-transformed mollifier is evaluated at all available wavenumbers (also outside the bandlimit) without additional truncation.
@@
where $\|\bm{a}\|_{2}=\sqrt{\sum_{j=1}^{n}|a_{j}|^{2}}$ denotes the standard $\ell_{2}$ vector norm. Absolute RMS errors are used when comparing numerical results to the theoretical error models in Section 5, which are formulated in terms of absolute errors. Relative $\ell_{2}$ errors are used when comparing different mollifier and window-function configurations, particularly when the relative error magnitude is important for the analysis.
@@ 7.3 Fast Ewald with PSWF split and PSWF window
When both the Ewald mollifier and the window function are based on PSWFs, the overall accuracy depends on two shape parameters: $c_{s}$, which controls the Ewald split, and $c_{w}$, which governs the window function. Proposition 2 shows that the aliasing error decays rapidly as $c_{w}$ increases, while Proposition 1 characterizes the truncation error associated with $c_{s}$. The total error may be viewed as approximately additive in these two contributions (see (67)) and is dominated by the larger of the two.
Figure 4: Absolute RMS errors for fast Ewald summation (Algorithm 2) using the PSWF split and PSWF window ($L=1$, $r_{c}=0.1$, $n=100$). Curves with symbols show numerical results, while black curves without symbols show theoretical error models. Dotted lines indicate the simple asymptotic exponential model $e^{-\{c_{s},c_{w}\}}$; solid lines correspond to the split-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{r_{c}}\,\|\bm{\rho}\|_{2}\,A_{s}c_{s}^{-1/2}e^{-c_{s}}$ (91); and dash–dotted lines correspond to the aliasing-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{L}\,\|\bm{\rho}\|_{2}\,A_{w}c_{w}^{1/2}e^{-c_{w}}$ (115). (a) Absolute RMS error vs. window support $P=(2/\pi)c_{w}$ for fixed $c_{s}$. (b) Absolute RMS error vs. number of Fourier modes $m$ (equivalently $c_{s}$) for fixed $P$.
@@
We observe that the error curves and plateaus in Figure 4 are well approximated by the error models (91) and (115). The theoretical predictions accurately capture both the slopes and the plateau levels, with only a slight overestimation of the error. This indicates that the chosen parameter values yield computed errors that are close to—and effectively bounded by—the prescribed tolerance $\varepsilon$.
Figure 5: Dependence of the absolute RMS error on the system size $n$ and the cutoff radius $r_{c}$ for fast Ewald summation with PSWF split and PSWF mollifier. (a) Window support vs. absolute RMS error for varying $n$. (b) Same data as in (a), rescaled by $\|\bm{\rho}\|_{2}$. (c) Window support vs. absolute RMS error for varying $r_{c}$. (d) Same data as in (c), rescaled by $\sqrt{r_{c}}$. Curves with symbols show numerical results, while black curves without symbols indicate theoretical error models: solid and dash-dotted lines correspond to the split-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{r_{c}}\,\|\bm{\rho}\|_{2}\,A_{s}c_{s}^{-1/2}e^{-c_{s}}$ (91), and the aliasing-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{L}\,\|\bm{\rho}\|_{2}\,A_{w}c_{w}^{1/2}e^{-c_{w}}$ (115) respectively.
@@
Figures 5(a)–(b) confirm that the absolute RMS error grows nearly linearly with $\|\bm{\rho}\|_{2}$. After rescaling by $\|\bm{\rho}\|_{2}$, the error curves nearly collapse, in agreement with the linearity of the potential. That this collapse is not exact is also expected: since the particle configurations are not perfectly uniformly random, the prefactor $C_{\rho}$ in Lemma 3 (see also Remark 10) need not be exactly one. In the worst-case scenario of strongly clustered particles, the dependence on $n$ may therefore be more pronounced. Nevertheless, the near-collapse observed here suggests that $C_{\rho}$ is close to one for the particle distributions considered in this example and that the proposed error models accurately capture the observed behavior.
@@ 8 Conclusions
We have presented a complete description of a new version of fast Ewald summation based on the first prolate spheroidal wave function of order zero (PSWF). While PSWFs have recently attracted attention in this context, our main contribution is a systematic and unified treatment: we define PSWF-based mollifier and window functions, establish their analytical properties, and integrate them into the fast Ewald framework.
The resulting PSWF-based split has several advantageous properties. Because PSWFs form an exact Fourier transform pair on a bandlimited interval, the Fourier-space representation is particularly simple, and the residual kernel becomes compactly supported, eliminating truncation errors in the real-space sum. Consequently, the total approximation error is governed entirely by Fourier truncation and aliasing. Owing to the superior spectral localization of PSWFs, a target accuracy can be reached with substantially fewer Fourier modes and smaller window supports than in Gaussian- and B-spline-based methods. Numerical experiments confirm these predictions and show significant reductions in computational cost at fixed accuracy.
We have also derived a detailed error analysis for both Fourier truncation and window-induced aliasing. The resulting sharp bounds and simplified closed-form error models exhibit exponential decay and enable transparent and reliable parameter selection, providing a practical way to balance accuracy and efficiency.
Future work will include extending the PSWF-Ewald method and the error analysis to other interaction kernels, such as those arising in Stokes flow and wave propagation, and investigating the potential benefits of oversampling to further accelerate fast Ewald summation.
@@ Appendix A PSWF asymptotics and approximations
In this section, we present the asymptotic behavior of $\psi_{0}^{c}$, along with corresponding approximations. These results are used to derive the closed-form approximations in Section 5.
@@
$1-\mu_{0}(c)=2\pi\int_{\mathbb{R}\setminus[-c,c]}|\widehat{\psi}_{0}^{c}(\omega)|^{2}\,d\omega=\|\widehat{\psi}_{0}^{c}\|^{2}_{L^{2}(\mathbb{R}\setminus[-c,c])},$ (131)
where $\mu_{0}(c)=c/(2\pi)\,\lambda_{0}^{2}$ is the first concentration eigenvalue associated with $\psi_{0}^{c}$ and $\|\psi_{0}^{c}\|_{L^{2}(-1,1)}=1$. Fuchs [13] proved the asymptotic behaviour
$1-\mu_{0}(c)\sim 4\sqrt{\pi c}\,e^{-2c}\qquad\text{as }c\to\infty,$ (132)
and therefore
$\|\widehat{\psi}_{0}^{c}\|_{L^{2}(\mathbb{R}\setminus[-c,c])}\sim C_{\mathrm{F}}\,c^{1/4}e^{-c},$ (133)
@@
For large $c$, the first PSWF eigenvalue of the differential operator satisfies [24]
@@
$\widehat{\psi}_{0}^{c}(c)=\lambda_{0}(c)\,\psi_{0}^{c}(1)$ (135)
shows that the decay of $\widehat{\psi}_{0}^{c}(c)$ determines the decay of $\psi_{0}^{c}(1)$. Using (134) and (135) gives
$\frac{\psi_{0}^{c}(1)}{\psi_{0}^{c}(0)}\sim\sqrt{\frac{c}{2\pi}}\,\frac{\widehat{\psi}_{0}^{c}(c)}{\psi_{0}^{c}(0)}.$ (136)
Motivated by the strong exponential localization of $\widehat{\psi}_{0}^{c}$ near $|\omega|=c$, we approximate
$\widehat{\psi}_{0}^{\,c}(c)\sim\|\widehat{\psi}_{0}^{c}\|_{L^{2}(\mathbb{R}\setminus[-c,c])}.$ (137)
@@
$\frac{\psi_{0}^{c}(1)}{\psi_{0}^{c}(0)}$ $\approx A_{2}\,c^{1/2}e^{-c},$ (142)
$\frac{2}{c\,\psi_{0}^{c}(0)}\sqrt{\frac{2\pi}{\lambda_{0}^{2}}-c}$ $\approx A_{3}\,c^{-1/2}e^{-c}.$ (143)
Here, (141) is obtained from (136) together with (133) and (138), and the expression for $\psi_{0}^{c}(1)$ follows from multiplying (142) by (140). The expression in (143) is fitted with an exponential, motivated by (132):
$\frac{2}{c\,\psi_{0}^{c}(0)}\sqrt{\frac{2\pi}{\lambda_{0}^{2}}-c}=\frac{2\sqrt{c}}{c\,\psi_{0}^{c}(0)\sqrt{\mu_{0}(c)}}\sqrt{1-\mu_{0}(c)}\sim C\,c^{-1/2}e^{-c}.$ (144)
@@
Figure 9: Curve-fitted approximations of key PSWF quantities; $E(c):=\frac{2}{c\,\psi_{0}^{c}(0)}\sqrt{\frac{2\pi}{\lambda_{0}^{2}}-c}$.
@@ B.1 Legendre expansion and eigenproblem
where only even Legendre polynomials appear due to symmetry. Here $P_{k}$ denotes the standard (unnormalized) Legendre polynomial; the normalized basis $\bar{P}_{k}(x)=\sqrt{k+\tfrac{1}{2}}\,P_{k}(x)$ used in [23] leads to an equivalent eigenproblem with rescaled coefficients. Substitution into the differential equation
$\frac{d}{dx}\Big((1-x^{2})\psi^{\prime}(x)\Big)+(\chi_{0}-c^{2}x^{2})\psi(x)=0$ (146)
leads to a three-term recurrence for the coefficients $\{a_{2k}\}$. After truncation at order $K$, a finite-dimensional symmetric tridiagonal eigenproblem is obtained. Its smallest eigenvalue corresponds to $\chi_{0}$ and the associated eigenvector provides the coefficients $a_{2k}$ [23, Theorem 10].
@@ B.2 Computation of eigenpairs
The tridiagonal structure permits efficient and accurate computation of the eigenvalue $\chi_{0}$ and the coefficient vector $\{a_{2k}\}$. In practice, Sturm bisection followed by inverse iteration is applied, as recommended in [23, Sec. 3.4]. This yields rapidly convergent approximations to $\psi_{0}^{c}$ with complexity $\mathcal{O}(K^{2})$.

### v2

@@ 6.2.1 Choosing $c_{s}$
Using the fitted model (89), we choose $c_{s}$ as
$c_{s}\approx\frac{1}{2}W\Big(2\,\big(B_{s}/\varepsilon\big)^{2}\Big),\qquad B_{s}:=\frac{1}{\sqrt{V}}A_{s}\|\bm{\rho}\|_{2}\sqrt{r_{c}},$ (120)
where $A_{s}\approx 5$. This choice provides accurate parameter selection across the practically relevant range of $c_{s}$ and yields close agreement with the observed absolute RMS errors.
A simpler, but less accurate alternative, is based on the asymptotic estimate [17, 19]
$c_{s}\approx\log(1/\varepsilon),$ (121)
which captures the correct exponential scaling and has the advantage of being parameter-free. It provides a reasonable estimate of $c_{s}$, although it may slightly overestimate the required bandwidth for small to moderate values of $c_{s}$ (see Figure 5).
In practice, both choices are effective.
@@ 6.3.1 Choosing $c_{w}$
To choose $c_{w}$, we assume that the split and window errors are balanced. Equating (114) and (89) yields
@@
Notably, the constants $A_{s}$ and $A_{w}$ share the same factors, so that their ratio is independent of $C_{\rho}$. In particular, we have
$\frac{A_{s}}{A_{w}}=\frac{4\sqrt{5}}{\pi^{3/2}}\approx 1.61.$ (127)
Solving (126) for $c_{w}$ using the Lambert $W$ function gives
$c_{w}\approx-\frac{1}{2}W\!\left(-2B_{w}^{2}\frac{e^{-2c_{s}}}{c_{s}}\right),\qquad B_{w}:=\frac{A_{s}}{A_{w}}\sqrt{\frac{r_{c}}{L}}.$ (128)
@@ 6.4 Parameter selection procedure
1: Solve $c_{s}=\frac{1}{2}W\big(2\,(B_{s}/\varepsilon)^{2}\big)$ for $c_{s}$, where $B_{s}:=\frac{1}{\sqrt{V}}A_{s}\|\bm{\rho}\|_{2}\sqrt{r_{c}}$; alternatively, use $c_{s}=\log(1/\varepsilon)$.
@@ 7.1 Experimental setup
Unless otherwise stated, we consider systems of $n=100$ particles uniformly distributed in $0,1)^{3}$, generated using MATLAB’s rand function. The particle charges $\rho_{j}$ are sampled independently from a standard normal distribution (mean 0, variance $1$) and subsequently shifted to enforce charge neutrality, $\sum_{j}\rho_{j}=0$. The choice $n=100$ is sufficient to capture the relevant convergence trends; larger systems exhibit the same slopes and offsets; see e.g. Figure 6. Unless otherwise stated, we set the cutoff radius to $r_{c}=0.1$ and the domain size to $L=1$, which provides reasonable Fourier-space resolution for testing purposes. For the split model, we use $A_{s}\approx 5$, and for the aliasing error model, $A_{w}\approx 3.1$, as discussed in Section 5.
We compute the approximate far-field potential values $\bm{\phi}^{\mathrm{far}}=(\phi^{\mathrm{far}}_{j})_{j=1}^{n}$ using Algorithm 2 and compare them with a reference solution $\bm{\phi}^{\mathrm{far}}_{\mathrm{ref}}=(\phi^{\mathrm{far}}_{\mathrm{ref},j})_{j=1}^{n}$, obtained by direct Ewald summation (Algorithm 1). In the reference calculation, we use a sufficiently large number of Fourier modes so that the truncation error is at least one order of magnitude below the target error level. The Fourier-transformed mollifier is evaluated at all available wavenumbers (also outside the bandlimit) without additional truncation.
@@
where $\|\bm{a}\|_{2}=\sqrt{\sum_{j=1}^{n}|a_{j}|^{2}}$ denotes the standard $\ell_{2}$ vector norm. Absolute RMS errors are used when comparing numerical results to the theoretical error models in Section 5, which are formulated in terms of absolute errors. Relative $\ell_{2}$ errors are used when comparing different mollifier and window-function configurations.
@@ 7.3 Fast Ewald with PSWF split and PSWF window
When both the mollifier and the window function are based on PSWFs, the overall accuracy depends on two shape parameters: $c_{s}$, which controls the kernel split, and $c_{w}$, which governs the window function. Proposition 2 shows that the aliasing error decays rapidly as $c_{w}$ increases, while Proposition 1 characterizes the truncation error associated with $c_{s}$. The total error may be viewed as approximately additive in these two contributions (see (72)) and is dominated by the larger of the two.
Figure 5: Absolute RMS errors for fast Ewald summation (Algorithm 2) using the PSWF split and PSWF window ($L=1$, $r_{c}=0.1$, $n=100$). Curves with symbols show numerical results, while black curves indicate theoretical error models. Dotted lines indicate the simple asymptotic exponential model $e^{-\{c_{s},c_{w}\}}$; solid lines correspond to the split-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{r_{c}}\,\|\bm{\rho}\|_{2}\,A_{s}c_{s}^{-1/2}e^{-c_{s}}$ (89) with $A_{s}\approx 5$; and dash–dotted lines correspond to the aliasing-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{L}\,\|\bm{\rho}\|_{2}\,A_{w}c_{w}^{1/2}e^{-c_{w}}$ (114) with $A_{w}\approx 3.1$. (a) Absolute RMS error vs. window support $P=(2/\pi)c_{w}$ for fixed $c_{s}$. (b) Absolute RMS error vs. number of Fourier modes $m$ (equivalently $c_{s}$) for fixed $P$.
@@
We observe that the error curves and plateaus in Figure 5 are well approximated by the error models (89) and (114). The theoretical predictions accurately capture both the slopes and the plateau levels. This indicates that the chosen parameter values yield computed errors that are close to the prescribed tolerance $\varepsilon$.
Figure 6: Dependence of the absolute RMS error on the system size $n$ and the cutoff radius $r_{c}$ for fast Ewald summation with PSWF split and PSWF mollifier. (a) Window support vs. absolute RMS error for varying $n$. (b) Same data as in (a), rescaled by $\|\bm{\rho}\|_{2}$. (c) Window support vs. absolute RMS error for varying $r_{c}$. (d) Same data as in (c), rescaled by $\sqrt{r_{c}}$. Curves with symbols show numerical results, while black curves indicate theoretical error models: solid lines correspond to the split-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{r_{c}}\,\|\bm{\rho}\|_{2}\,A_{s}c_{s}^{-1/2}e^{-c_{s}}$ (89), and dash–dotted lines to the aliasing-error model $\varepsilon=\tfrac{1}{\sqrt{V}}\sqrt{L}\,\|\bm{\rho}\|_{2}\,A_{w}c_{w}^{1/2}e^{-c_{w}}$ (114).
@@
Figures 6(a)–(b) confirm that the absolute RMS error grows linearly with $\|\bm{\rho}\|_{2}$. After rescaling by $\|\bm{\rho}\|_{2}$, the error curves collapse, in agreement with the linearity of the potential.
@@ 8 Conclusions
We have presented a complete description of fast Ewald summation based on the first prolate spheroidal wave function of order zero (PSWF). While PSWFs have recently attracted attention in this context, our main contribution is a systematic and unified treatment: we define PSWF-based mollifier and window functions, establish their analytical properties, and integrate them consistently into the fast Ewald framework.
The resulting PSWF-based split has several advantageous properties. Because PSWFs form an exact Fourier transform pair on a bandlimited interval, the Fourier-space representation is particularly simple, and the residual kernel becomes compactly supported, eliminating truncation errors in the real-space sum. Consequently, the total approximation error is governed by Fourier truncation and aliasing. Owing to the superior spectral localization of PSWFs, a prescribed accuracy can be reached with substantially fewer Fourier modes and smaller window supports than in Gaussian- and B-spline-based methods. Numerical experiments confirm these predictions and demonstrate significant reductions in computational cost at fixed accuracy.
We have derived a detailed error analysis for both Fourier truncation and window-induced aliasing. The resulting error estimates and simplified closed-form models exhibit exponential decay and enable transparent and reliable parameter selection, providing a practical framework for balancing accuracy and computational cost.
Future work includes extending the PSWF-Ewald method and the error analysis to other interaction kernels, such as those arising in Stokes flow and wave propagation, and investigating the potential benefits of oversampling to further accelerate fast Ewald summation.
@@ Appendix A PSWF asymptotics and approximations
In this section, we summarize the asymptotic behavior of $\psi_{0}^{c}$, along with corresponding approximations. These results are used to derive the closed-form approximations in Section 5.
Define
[…]
@@
$1-\mu_{0}(c)=2\pi\int_{\mathbb{R}\setminus[-c,c]}|\hat{f}(\omega)|^{2}\,d\omega=2\pi\|\hat{f}\|^{2}_{L^{2}(\mathbb{R}\setminus[-c,c])},$ (134)
where $\mu_{0}(c):=c/(2\pi)\,\lambda_{0}^{2}$ denotes the first concentration eigenvalue associated with $\psi_{0}^{c}$ and where $\|f\|_{L^{2}(-1,1)}=\|\psi_{0}^{c}\|_{L^{2}(-1,1)}=1$. Fuchs [13] proved the asymptotic behaviour
$1-\mu_{0}(c)\sim 4\sqrt{\pi c}\,e^{-2c}\qquad\text{as }c\to\infty.$ (135)
Taking the square root yields
$\|\hat{f}\|_{L^{2}(\mathbb{R}\setminus[-c,c])}\sim C_{\mathrm{F}}\,c^{1/4}e^{-c},$ (136)
@@
For large $c$, the first PSWF eigenvalue of the integral eigenvalue equation (23) satisfies [23]
@@
$\hat{f}(c)=\lambda_{0}(c)\,\psi_{0}^{c}(1)$ (138)
shows that the decay of $\hat{f}(c)$ determines the decay of $\psi_{0}^{c}(1)$. Using (137) and (138) gives
$\frac{\psi_{0}^{c}(1)}{\psi_{0}^{c}(0)}\sim\sqrt{\frac{c}{2\pi}}\,\frac{\hat{f}(c)}{\psi_{0}^{c}(0)}.$ (139)
Motivated by the strong exponential localization of $\hat{f}$ near $|\omega|=c$, we assume that the pointwise value $\hat{f}(c)$ is of the same order as the tail norm, and write
$\hat{f}(c)\approx\|\hat{f}\|_{L^{2}(\mathbb{R}\setminus[-c,c])}.$ (140)
@@
$\frac{\psi_{0}^{c}(1)}{\psi_{0}^{c}(0)}$ $\approx A_{2}\,c^{1/2}e^{-c},.$ (145)
Here, (144) is obtained from (139) together with (136) and (141), and (145) follows by dividing (144) by (143).
@@
Figure 10: Curve-fitted approximations of key PSWF quantities.
@@ B.1 Legendre expansion and eigenproblem
where only even Legendre polynomials appear due to symmetry. Here $P_{k}$ denotes the standard (unnormalized) Legendre polynomial; the normalized basis $\bar{P}_{k}(x)=\sqrt{k+\tfrac{1}{2}}\,P_{k}(x)$ used in [24] leads to an equivalent eigenproblem with rescaled coefficients. Substitution into the differential equation (30) leads to a three-term recurrence for the coefficients $\{a_{2k}\}$. After truncation at order $K$, a finite-dimensional symmetric tridiagonal eigenproblem is obtained. Its smallest eigenvalue corresponds to $\chi_{0}$ and the associated eigenvector provides the coefficients $a_{2k}$ [24, Theorem 10].
@@ B.2 Computation of eigenpairs
The tridiagonal structure permits efficient and accurate computation of the eigenvalue $\chi_{0}$ and the coefficient vector $\{a_{2k}\}$. In practice, Sturm bisection followed by inverse iteration is applied, as recommended in [24, Sec. 3.4]. This yields rapidly convergent approximations to $\psi_{0}^{c}$ with complexity $\mathcal{O}(K^{2})$.